OpenStudy (anonymous):

solve the given equation 2 cos^2 θ − cos θ − 1 = 0

OpenStudy (anonymous):

factor

OpenStudy (imstuck):

yep...factor just as you would a "regular" polynomial of second degree. It works beautifully. If you need help, we can guide you...

OpenStudy (imstuck):

Factor it into parenthesis like this: (2 cos theta + )(cos theta - ). Fill in the blanks. Do you know how to do this?

OpenStudy (anonymous):

I've got \[2\pi/3+2\pi k\]and \[4\pi/3+2\pi k\] is it correct?

OpenStudy (anonymous):

that's for theta

OpenStudy (midhun.madhu1987):